OpenStudy (anonymous):

sqrt(6-x)-2 / sqrt(3-x)-1 find the limit as x->2

myininaya (myininaya):

satellite go ahead and tell him that he might have to multiply bottom and top by the conjugate of both the top and the bottom

OpenStudy (anonymous):

i did that already, but it was deleted!

OpenStudy (anonymous):

as i was writing it, actually as i was trying to post it, it vanished!

OpenStudy (anonymous):

so you go ahead because i got annoyed

myininaya (myininaya):

\[\lim_{x \rightarrow ?}\frac{\sqrt{6-x}-2}{\sqrt{3-x}-1} \cdot \frac{\sqrt{6-x}+2}{\sqrt{3-x}+1} \cdot \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\]

myininaya (myininaya):

\[=\lim_{x \rightarrow 2}\frac{(6-x)-2}{(3-x)-1} \cdot \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\]

myininaya (myininaya):

\[=\lim_{x \rightarrow 2}\frac{-x+4}{-x+2} \cdot \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\] maybe i should had tried just multiplying top and bottom by the conjugate of the bottom

OpenStudy (anonymous):

no you made an arithmetic error

myininaya (myininaya):

oh i see -2(2)=-4

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

damn site is so slow it is not really worth the wait tonight. i post, and then noting happens but the zeros will cancel because your numerator should be \[2-x\]

myininaya (myininaya):

\[\lim_{x \rightarrow ?}\frac{\sqrt{6-x}-2}{\sqrt{3-x}-1} \cdot \frac{\sqrt{6-x}+2}{\sqrt{3-x}+1} \cdot \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2} \] \[=\lim_{x \rightarrow 2}\frac{(6-x)-4}{(3-x)-1} \cdot \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\]

myininaya (myininaya):

\[=\lim_{x \rightarrow 2}\frac{-x+2}{-x+2} \cdot \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\]

OpenStudy (anonymous):

\[=\lim_{x \rightarrow 2}\frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\]

myininaya (myininaya):

\[=\lim_{x \rightarrow 2} \frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}=\frac{\sqrt{3-2}+1}{\sqrt{6-2}+2}\]

OpenStudy (anonymous):

like this in the morning too. hope it is not a trend

OpenStudy (anonymous):

tyvm

myininaya (myininaya):

this site is going freakin' slow tonight :(

OpenStudy (anonymous):

i knw